OpenStudy (anonymous):

FAN AND MEDAL! I need help please? (:

OpenStudy (anonymous):

SOME OF THEM I THINK THAT I KNOW...BUT IM NOT SURE..I REALLY NEED HELP GUYS. PLEASE HELP? I DON'T UNDERSTAND ANY OF THIS :'( A school puts on a play. The play costs $1,200 in expenses. The students charge $4.00 for tickets. There will be one performance of the play in an auditorium that seats 500 people. What is the domain of the function that shows the profit as a function of the number of tickets sold? The domain is all real numbers from negative 1,200 to positive 2,000 The domain is all integers from negative 1,200 to positive 2,000. The domain is the integers from 0 to 500. The domain is all real numbers from 0 to 500. ............................................................................................................................... Question 2 The graph shows the temperature of ice cream in an ice chest during an overnight trip. What is the domain of this function? Graph showing a line segment that goes from ordered pairs 0, negative 10 to 30, 30. The domain is negative 10 through 30. The domain is all integers from negative 10 through 30. The domain is all real numbers 0 through 30. The domain is all integers from 0 through 30. .............................................................................................................................. Question 3 Your friend has $80 when he goes to the fair. He spends $4 to enter the fair and $12 on food. Rides at the fair cost $1.25 per ride. Which function can be used to determine how much money he has left over after x rides? f(x) = 1.25x + 64 f(x) = -16x + $80 f(x) = -1.25x + 64 f(x) = -1.25x - 64 .............................................................................................................................. Question 4 Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = ______ Graph of f of x and g of x. f of x equals 1 over 3 x minus 2 and g of x equals 1 over 3 x plus 3. Numerical Answers Expected! Answer for Blank 1: ............................................................................................................................. Question 5 A student is running a 3-kilometer race. He runs 1 kilometer every 2 minutes. Select the function that describes his distance from the finish line. f(x) = - 1 over 3x + 2 f(x) = - 1 over 2x + 3 f(x) = 1 over 2x + 3 f(x) = 1 over 3x + 2 ........................................................................................................................... Question 6 A coal car on a train weighs 30 tons plus 1 ton per cubic yard of coal x that it carries. The total weight of a coal car is: f(x) = x + 30. How will the graph of this function change if the coal car weight is changed to 26 tons? The line will shift vertically up by 4 tons. The line will shift vertically up 26 tons. The line will shift vertically down by 4 tons. The line will shift vertically down by 26 tons. ............................................................................................................................. Question 7 A semiconductor manufacturing company employs 1,000 people. The average weekly salary of each employee is $750. Select the graph that correctly shows the total cost to the company of paying its employees during a year. Graph of Weeks versus Dollars in millions. Line segment between points zero comma zero and 52 comma 39 million. Graph of Weeks versus Dollars in millions. Ray starting at point zero comma zero and passing through point 52 comma 39 million. Graph of Dollars in millions versus weeks. Ray starting at point zero comma zero and passing through point 39 million comma 52. Graph of Dollars in millions versus weeks. Line segment between points zero comma zero and 39 million comma 52. ............................................................................................................................... Question 8 A cell phone plan has a monthly cost that is shown in the table below. What is the correct statement regarding the average rate of change during the 40-minute time of talk? Total minutes of talk time Monthly cost of cell phone 0 $19.95 10 $20.45 20 $20.95 30 $21.45 40 $21.95 The average rate of change is $0.50, meaning that for each minute of talk time, the monthly bill increases by $0.50. The average rate of change is $0.50, meaning that for every ten minutes of talk time, the bill increases by $0.50. The average rate of change is $0.05, meaning that for each minute of talk time, the monthly bill increases by $0.05. The average rate of change is $0.05, meaning that for every ten minutes of talk time, the bill increases by $0.05. ............................................................................................................................ Question 9 Two sisters decide to take a series of acting lessons. One decides to pay a flat rate of $75, shown as f(x). The other just wants to pay $5 per lesson, shown as g(x). Which function shows the correct combination of these two functions to represent the amount the parent should pay, shown as h(x)? f(x) = 75, g(x) = 5x, h(x) = 5x + 75 f(x) = 75x, g(x) = 5, h(x) = 75x + 5 f(x) = 75, g(x) = 5, h(x) = 5 + 75 f(x) = 75x, g(x) = 5x, h(x) = 75x + 5x

OpenStudy (anonymous):

@hba @jigglypuff314 @iPwnBunnies @AngelWilliams16 @stabar @Somy @SiennaSurfer @it_meh_miranda @makayla_finkbeiner_ @m&msdoodle @Secret-Ninja @Passionateboyy14 @melody16 @MicroBot @hugsnotughs @lovelyharmonics @JoMary @lxl-rhea-lxl

OpenStudy (anonymous):

Is this a test?

OpenStudy (anonymous):

No, my tests are MUCH longer! This is an assignment I have to do and it is actually 20 questions long but I know the answers to the rest. It's just these few are KILLING me! :(

jigglypuff314 (jigglypuff314):

alright then for the first one Cost = (price per ticket) times (number of tickets) if you compare it to how Domain and Range works where Domain = x values and Range = y values C = (p)*(t) y = (a)*(x) ^ domain refers to number of tickets so how many tickets are you able to sell?

OpenStudy (anonymous):

Would it be A? All real numbers from negative 1,200 to positive 2,000?

jigglypuff314 (jigglypuff314):

not quite :) hint: number of tickets = number of people how many people can you possibly have?

OpenStudy (anonymous):

500

OpenStudy (anonymous):

Oh...I looked at it wrong. We are comparing C = (p)*(t) and y = (a)*(x) So t is similiar to x, right?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

So D? :) ..please say yes!

jigglypuff314 (jigglypuff314):

hmmm well all real numbers includes fractions right? can you have half a person? O_O

OpenStudy (anonymous):

Okay...your right! I forgot that all real numbers consists of fractions...So then it would be integers NOT all real numbers

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

Awesome! So what about number 2? It is a graph?

OpenStudy (anonymous):

Domain is x values...so, it would be B? I think?

OpenStudy (anonymous):

No, not B...Because the -10 is a y value...so...

OpenStudy (anonymous):

C again? Since with this graph, you CAN have fractions?

jigglypuff314 (jigglypuff314):

That is correct :) Good Job! :D

OpenStudy (anonymous):

Awesome! Okay...next one? (: And thank you for the help!

OpenStudy (anonymous):

I think it's D...but I'm not really sure. I feel like that is wrong...

jigglypuff314 (jigglypuff314):

it's close though :) you start with a Positive 80 then subtract 4 for the entrance fee then subtract 12 for the food do you get a Positive or Negative number?

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

So it's C...AGAIN! lol I see where I messed up

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

Okay...Now, the next one is the hardest for me...It is a fill in the blank too (if that matters).

jigglypuff314 (jigglypuff314):

I'm not sure if it is 1/(3x) or (1/3)x but it doesn't matter either way I suppose... turning what it gave into math-speak ;) f(x) = 1/3x - 2 g(x) = 1/3x + 3 and they gave g(x) = f(x) + k so looking between g(x) and f(x), how do you think you could get from f(x) to g(x) ?

OpenStudy (anonymous):

I guess you could say: 1/3(1/3+3) - 2

OpenStudy (anonymous):

You substitute the x in f(x) for the function g(x) right?

jigglypuff314 (jigglypuff314):

ummmmm try g(x) = f(x) + k -f(x) -f(x) k = g(x) - f(x)

OpenStudy (anonymous):

But where does the number come in? Because it is asking for a numerical answer?

jigglypuff314 (jigglypuff314):

k = g(x) - f(x) substitute in f(x) = 1/3x - 2 g(x) = 1/3x + 3

OpenStudy (anonymous):

So, K= (1/3x+3) - (1/3x-2) ?

jigglypuff314 (jigglypuff314):

yes :) then simplify

OpenStudy (anonymous):

5?

jigglypuff314 (jigglypuff314):

Correct ^_^

OpenStudy (anonymous):

Awesome! OMG Thank you a lot! You are REALLY helping! Only a few more (:

OpenStudy (anonymous):

For the next one, I think it is D. I asked somebody else earlier and they agreed. They said it is D because its a 3 kilo race and it 1 kilometer per 2 minutes. So, it would be 1 kilo divided by 2 minutes plus the total number of kilometers which is 3...Is that right?

OpenStudy (anonymous):

I mean C

jigglypuff314 (jigglypuff314):

close :P as time goes on, he gets closer to the finish line right? so (total distance) - (distance already ran) = (distance left) where (distance already ran) = (rate)*(time)

OpenStudy (anonymous):

So, the total distance is 3 the distance already ran is 1 and the distance left is 2. So, since there is no choice for 3 over 1 = 2, would I make it negative?

jigglypuff314 (jigglypuff314):

rate = (1km in 2 min) = 1/2 total distance = 3 time = x

OpenStudy (anonymous):

I'm confused...I thought you said it wouldn't be 1/2x +3 ?

jigglypuff314 (jigglypuff314):

(total distance) - (distance already ran) = (distance left) ^ there's a negative somewhere because 1/2x + 3 means that you're running away from the finish line xD

OpenStudy (anonymous):

Okay...So I need to do the opposite, which would be -1/2x +3

jigglypuff314 (jigglypuff314):

that's correct :)

OpenStudy (anonymous):

Awesome. For the next one, I think I have the answer. The line shifts vertically down by 4 tons

jigglypuff314 (jigglypuff314):

that is correct :)

OpenStudy (anonymous):

Great! I have the last 2..I just need the next one about the semiconductors

jigglypuff314 (jigglypuff314):

time is always the independent variable so in this case "Weeks" should go along the x axis